# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

### v1

@@ 3.1.1 Robust Graph Learning
(2) LLM-enhanced methods Recently, researchers have begun exploring the deep integration of semantic reasoning and incomplete graph learning. Methods based on LLMs [76, 75, 77, 79, 80, 78] have pioneered a new approach of “semantic completion–topology optimization” co-evolution, leveraging the strong generative and reasoning capabilities of LLMs.
For instance, Sun et al. [76] proposes an LLM-based architecture that effectively tackles graph incompleteness by leveraging semantic reasoning to infer missing edges and nodes. Specifically, LLMs can identify and establish plausible connections between nodes based on their semantic similarity, thereby mitigating the common challenge of missing or incomplete relationships in real-world graphs. Experimental results indicate that these LLM-enhanced methods surpass traditional approaches—such as basic link prediction or imputation techniques—by producing more accurate and semantically meaningful graph structures. Similarly, LLM4NG [75] employs LLMs to generate contextually relevant nodes, enriching sparse or incomplete graph data, particularly in few-shot learning scenarios. This approach significantly improves graph completeness by addressing gaps that conventional methods (e.g., imputation) often fail to resolve effectively.
Meanwhile, SPLLM [77] tackles the challenge of incompleteness in high-dimensional sensor graphs, where missing values disrupt both spatial and temporal dependencies. It integrates a spatiotemporal fusion GCN with LLM-based fine-tuning, using external domain knowledge (e.g., road networks, weather data) to infer missing patterns and complete partial features more effectively. UnIMP [81], in a similar spirit, targets mixed-type tabular data by leveraging LLMs with bidirectional high-order message passing on cell-oriented hypergraphs. This approach addresses incompleteness arising from semantic and structural heterogeneity, which traditional imputation methods struggle with. Figure 5 illustrates how LLMs contribute to graph completion by extracting domain-specific semantic features from textual descriptions or external knowledge. These features serve as compensatory signals for missing nodes, edges, or attributes in sparse or incomplete graphs.
@@
To further close the semantic gap introduced by graph incompleteness, several works propose end-to-end frameworks combining LLMs and graph models. For instance, GraphLLM [79] introduces a unified architecture where a graph encoder maps node features into the LLM’s semantic space, enabling it to reason over incomplete or noisy node attributes using attention-gated fusion of structure and context. This helps, for example, in citation networks or recommendation systems where node metadata may be sparse or outdated. PROLINK [80] enhances inductive reasoning on low-resource knowledge graphs by generating structural prompts and using LLMs to fill in gaps from limited graph and text data — particularly valuable in biomedical or emerging domains with sparse curation. Lastly, LLM-GNN [78] proposes a semi-supervised approach where LLMs annotate a small subset of nodes, helping GNNs generalize from minimal labeled data — addressing the classic label incompleteness challenge in large-scale graphs.
Empirical results in the literature indicate that LLM-augmented models achieve higher accuracy on incomplete graph benchmarks, especially when missingness is significant.
@@
Figure 5: The framework addresses incomplete graph structures by integrating structural signals from GNNs and semantic priors from LLMs. The “Align / Fusion / Distillation” module enables the model to compensate for missing graph information by leveraging textual context, resulting in enriched and unified node attribute/edge/label representations
@@ 3.1.2 Few-shot Graph Learning
(1) Traditional Methods for Few-Shot Learning FSL aims to enable effective model generalization from a limited number of labeled samples. In graph domains, FSL not only addresses label scarcity but also tackles structural challenges such as missing links and infrequent or long-tail relations [119, 120, 121, 122, 123]. To overcome these limitations, traditional approaches focus on explicitly modeling graph structures and relational semantics under low-resource conditions.
For example, Gmatching [119] pioneered the use of single-hop neighborhood structures and relation prototype matching to handle rare relation prediction. GFL [120] establishes a transferable metric space by projecting auxiliary knowledge onto the target graph. FSRL [121] proposes a heterogeneous relation meta-learning framework with a reference representation generator to decouple relation semantics. FAAN [122] introduces an adaptive encoder and aggregator to model entity-reference contributions, significantly enhancing few-shot knowledge graph completion through fine-grained semantic representations. For real-world applications, MetaHG [123] addresses drug trafficking detection on Instagram by constructing a heterogeneous graph and optimizing node representations via meta-learning, outperforming existing methods.
@@
Despite these advancements, existing methods often rely on hand-crafted architecture designs and still face inefficiencies in handling complex graph structures, along with limited capabilities for structured and interpretable reasoning.
(2) Few-Shot Learning in LLMs Few-shot learning within LLMs has seen remarkable progress, particularly in tasks such as graph data analysis [84], link prediction [82], anomaly detection [83], table understanding [85], and knowledge graph question answering (KGQA) [86, 87].
LLMDGCN [84] introduces a degree-aware prompt tuning framework, encoding graph structures as positional embeddings for LLMs. This method not only filled missing node labels but also restored missing edges through an edge prediction module, improving classification accuracy and robustness in low-label scenarios. For link prediction in graph data, LinkGPT [82] extends LLMs’ few-shot learning capabilities by incorporating instruction tuning and retrieval-based re-ranking mechanisms. This approach significantly improved model reasoning under limited training data, yielding substantial performance gains on real-world datasets. In anomaly detection for dynamic graphs, AnomalyLLM [83] combines LLM pretraining knowledge with few-shot learning strategies, using a dynamic-aware encoder to generate edge representations and incorporating word embedding prototypes for edge reprogramming, successfully enhancing few-shot anomaly detection effectiveness. For table understanding tasks, the HGT framework [85] utilizes heterogeneous graph-enhanced LLMs, integrating soft prompts and instruction tuning to effectively align table semantics with LLM parameter knowledge, leading to more efficient table parsing in few-shot scenarios. In KGQA, FlexKBQA [86] improves question-answering performance in low-resource settings by converting knowledge base programs into natural language questions, enabling synthetic data generation. Meanwhile, the zero-shot KGQA framework [87] combines LLMs with smaller models, significantly reducing generation costs while achieving performance comparable to fully supervised models.
@@
In summary, while few-shot graph learning remains challenging, LLMs act as few-shot learners by nature – they can leverage their pre-trained knowledge to make sense of novel tasks with limited data. Results show that even with minimal graph information, an LLM can propose connections or classifications that align with human knowledge, essentially providing a powerful prior to the graph model.
@@ 3.1.3 Knowledge Graph Completion
Early studies focused on converting knowledge graph triples into textual sequences to validate the feasibility of LLM-based generation for KGC. FSKG [92] introduces an LLM-based generation and refinement framework to address data scarcity in long-tail scenarios. By employing multi-stage knowledge distillation, MuKDC effectively generates supplementary knowledge, improving KGC performance. KGs-LLM [91] proposes an innovative knowledge graph generation method utilizing generative LLMs (e.g., GPT-3.5). By leveraging zero-shot prompting and a strategy that requires no external knowledge, this approach iteratively extracts graph components without manual annotation, making it highly scalable and effective in domain-specific datasets. Similarly, KGLLM [93] transforms knowledge graph triples into text sequences and leverages LLMs for prediction, achieving state-of-the-art performance in triple classification and relation prediction. Compared to traditional methods, this approach enables efficient fine-tuning of smaller models, particularly lightweight LLMs such as LLaMA-7B. To mitigate the long-tail entity issue in KGC, KICGPT [94] integrates an LLM with a triple-based KGC retriever, encoding structural information into prompts using contextual learning strategies. This method requires no additional training costs and performs exceptionally well in few-shot settings.
To address the structural information loss inherent in text-based methods, researchers have developed multi-level prompt encoding techniques. RL-LLM [95] applies few-shot learning and multi-prompt optimization for relation prediction in e-commerce knowledge graphs, achieving performance comparable to human-labeled data in some cases, showcasing the potential of LLMs in such tasks. GoG [96] introduces a ”think-search-generate” approach, enabling the generation of new triples without additional training, thereby improving incomplete knowledge graph reasoning. This framework effectively integrates LLMs with knowledge graphs, significantly enhancing question-answering over incomplete KGs. LLM-KGC [88] embeds ontology knowledge and graph structure into LLM prompts, combining topological information with contextual learning techniques to achieve effective KGC across different settings. Experiments show that this approach outperforms baseline models across multiple datasets, particularly excelling in inductive reasoning tasks where strong inference capabilities are required. KoPA [97] introduces a knowledge prefix adapter, incorporating structural information into the LLM inference process. By converting graph entities and relations into textual representations suitable for LLM input, KoPA significantly enhances KGC performance, strengthening structure-aware reasoning in KGC tasks.
Building upon the advancements in LLM-driven KGC, several frameworks have further optimized few-shot KGC (FKGC), entity alignment, prompt engineering, and structural integration to enhance performance and scalability. MuKDC [101] employs multi-level knowledge distillation to generate supplementary knowledge while maintaining coherence and reliability, excelling in long-tail FKGC tasks through multimodal knowledge generation and consistency evaluation. LLMKG [98] enhances zero-shot and one-shot KGC using Mixtral-8x7b and GPT-4, demonstrating that well-designed prompts allow LLMs to perform effectively under various evaluation metrics. DIFT [99] improves entity alignment by integrating lightweight models and discriminative prompts, preventing alignment errors while significantly boosting KGC results. CP-KGC [100] enhances semantic richness in quantized LLMs through contextually adaptive prompts, effectively handling polysemous entities and pushing the boundaries of text-based KGC. Beyond textual enhancements, structural integration methods such as GLTW [90] and GS-KGC [89] further refine KGC. GLTW improves local and global structure encoding by integrating Graph Transformers with LLMs, while GS-KGC enhances contextual reasoning by leveraging subgraph-based question-answering and negative sampling, significantly improving LLM accuracy in missing triple identification.
@@ 3.2 Imbalance in Graphs
This section reviews the application of traditional methods and LLMs for learning from imbalanced graph data. The literature is categorized into two key research areas: Class Imbalance and Structure Imbalance. The relevant references and categorization are presented in Table 2.
Table 2: Literature Summary of LLM-based Methods for Handling Imbalance in Graphs.
Domain Task Reference (Method—Citation) Class Imbalance Graph Learning Node Classification LLM4NG[124], LLM-GNN[78], G2P2[125] LA-TAG[126], GSS-Net[127], TAGrader[128] LLM-SBCL[129], SEGA[130], SocioHyperNet[131] Cella[132], LLM-TIKG[133], ANLM-assInNNER[134] LLM-HetGDT[135] Prediction LKPNR[136], LLM-DDA[137] Graph Completion KICGPT[94], KGCD[138] Foundation Model GraphCLIP[139] Structure Imbalance Graph Learning Node Classification GraphEdit[140] Graph Completion SATKGC[141], MPIKGC[142], LLM4RGNN[143]
@@ 3.2.1 Class-Imbalanced Graph Learning
Another approach focuses on the contextual understanding and reasoning capabilities of LLMs, mining implicit relationships of minority classes through structure-aware reasoning. Both LLM-GNN [78] and Cella [132] proposed a zero-shot classification framework based on LLM semantic reasoning. By parsing node attributes and neighbor context, they directly generate category prediction confidence and get rid of the dependence on labeled data. G2P2 [125] introduces a graph-guided pre-training prompt framework, using LLMs to generate augmented text related to graph structure, enhancing the discriminability of low-resource classes in heterogeneous spaces. Meanwhile, LLM-TIKG [133] builds a threat intelligence knowledge graph, using LLMs to extract attack pattern association rules from unstructured text, and enhancing the representation of minority class threat entities through heterogeneous graph attention.
Other methods focus on external knowledge injection and transfer, mining long-tail information from open domains or domain-specific knowledge to improve the generalization ability of the model. KICGPT [94] dynamically builds prompt templates aligned with the graph structure based on contextual knowledge injection, enhances the reasoning ability of long-tail relations, and significantly improves the knowledge graph completion performance of sparse categories. LKPNR [136] combines knowledge graphs with LLM user preference reasoning to solve the cold start problem in long-tail news recommendation through semantic path enhancement. KGCD [138] uses LLM to generate pseudo-triplets based on logical rules to guide the knowledge graph completion model to focus on long-tail relations. At the same time, the method in [128] proposes a hierarchical knowledge distillation framework to transfer the open-domain semantic knowledge of LLM to the graph encoder to improve the generalization ability in the scenario of uneven categories. GraphCLIP [139] uses the cross-modal mapping capability of LLM to construct a graph-text comparison alignment framework to alleviate the problem of uneven categories in the target domain. Finally, LLM-DDA [137] integrates the disease-drug mechanism representation obtained by LLM reasoning into graph reconstruction in the biomedical scenario, and improves the accuracy of drug repurposing prediction for rare diseases through a multimodal attention mechanism.

### v2

@@ 3.1.1 Robust Graph Learning
(2) LLM-enhanced methods Recently, researchers have begun exploring the deep integration of semantic reasoning and incomplete graph learning. Methods based on LLMs [84, 41, 87, 91, 93, 89] pioneer a new paradigm that combines semantic completion and structural refinement, leveraging the strong generative and reasoning capabilities of LLMs.
Figure 5 illustrates how LLMs can extract domain-specific knowledge and use it to compensate for missing nodes, edges, or attributes. For example, Sun et al. [84] demonstrate that LLMs can infer missing nodes and edges by recognizing semantic similarities between entities. In this way, the model does not just “guess” a connection statistically, but proposes links that are meaningful in the real world. Building on this idea, LLM4NG [41] generates new, contextually relevant nodes, which is especially helpful in few-shot settings where the graph is very sparse. Both methods show that LLMs can enrich incomplete graphs more effectively than traditional imputation. Meanwhile, SPLLM [87] focuses on traffic sensor networks, where missing values break both spatial and temporal dependencies. By combining spatiotemporal GCNs with LLM fine-tuning, and incorporating external knowledge such as road maps or weather data, it achieves more reliable predictions. Similarly, UnIMP [97] works on tabular data by treating it as a hypergraph and using LLMs to fill in missing values, even when the data is semantically or structurally heterogeneous.
@@
To further close the semantic gap introduced by graph incompleteness, several works propose end-to-end frameworks combining LLMs and graph models. For instance, GraphLLM [91] introduces a unified architecture where a graph encoder maps node features into the LLM’s semantic space, enabling it to reason over incomplete or noisy node attributes using attention-gated fusion of structure and context. This helps, for example, in citation networks or recommendation systems where node metadata may be sparse or outdated. PROLINK [93] enhances inductive reasoning on low-resource knowledge graphs by generating structural prompts and using LLMs to fill in gaps from limited graph and text data — particularly valuable in biomedical or emerging domains with sparse curation. Lastly, LLMGNN [89] proposes a semi-supervised approach where LLMs annotate a small subset of nodes, helping GNNs generalize from minimal labeled data — addressing the classic label incompleteness challenge in large-scale graphs.
@@
Figure 5: The framework addresses incomplete graph structures by integrating structural signals from GNNs and semantic priors from LLMs. The “Align / Fusion / Distillation” module enables the model to compensate for missing graph information by leveraging textual context, resulting in enriched and unified node attribute/edge/label representations.
@@ 3.1.2 Few-shot Graph Learning
(1) Traditional Methods for Few-Shot Learning
FSL in graph domains aims to generalize effectively from limited labeled samples, addressing not only label scarcity but also structural issues such as missing links and long-tail relations [160, 161, 162, 163, 164]. Early approaches primarily focused on explicitly modeling local graph structures and relational semantics under low-resource settings. For instance, Gmatching [160] pioneered the integration of single-hop neighborhood information with relation prototype matching, providing a foundation for rare relation prediction. Building on this, GFL [161] established a transferable metric space by projecting auxiliary knowledge into the target domain, improving cross-graph adaptability.
Subsequent works moved toward more sophisticated meta-learning strategies. FSRL [162] introduced a heterogeneous relation meta-learning framework that decouples relation semantics through a reference representation generator, enabling better handling of heterogeneous relation types. Complementing this semantic focus, FAAN [163] designed an adaptive encoder–aggregator mechanism to dynamically weight entity-reference contributions, thereby producing fine-grained semantic representations that significantly enhance few-shot knowledge graph completion. Extending FSL techniques to real-world applications, MetaHG [164] demonstrated the utility of meta-learning in heterogeneous environments by detecting drug trafficking on Instagram.
@@
Despite these advances, traditional FSL methods remain constrained by hand-crafted architecture designs, often lacking efficiency when dealing with large, complex graph structures and exhibiting limited capability for structured and interpretable reasoning.
(2) Few-Shot Learning in LLMs Few-shot learning has emerged as a natural strength of LLMs, enabling them to handle graph-related tasks with limited labeled data by leveraging rich prior knowledge from pretraining. Recent work demonstrates their effectiveness in diverse scenarios such as graph classification [106], link prediction [99], anomaly detection [102], table understanding [108], and knowledge graph question answering (KGQA) [114, 118]. For instance, LLMDGCN [106] incorporates degree-aware prompt tuning with graph-encoded positional embeddings, enabling both node classification and edge recovery in low-label settings. In link prediction, LinkGPT [99] combines instruction tuning with retrieval-based re-ranking to enhance reasoning under sparse supervision. AnomalyLLM [102] integrates dynamic-aware encoding and prototype-based edge reprogramming for improved detection in dynamic graphs. Beyond pure graph tasks, HeGTa [108] aligns table semantics with LLM knowledge via heterogeneous graph-enhanced soft prompts, while FlexKBQA [114] and a zero-shot KGQA approach [118] reformulate KB queries into natural language to generate synthetic training data or reduce reliance on full supervision. Collectively, these results indicate that LLMs can generalize to novel graph tasks with minimal examples, acting as strong priors that bridge missing or noisy information.
@@
In summary, while few-shot graph learning remains challenging, LLMs act as few-shot learners by nature – they can leverage their pretrained knowledge to make sense of novel tasks with limited data. Results show that even with minimal graph information, an LLM can propose connections or classifications that align with human knowledge, essentially providing a powerful prior to the graph model.
@@ 3.1.3 Knowledge Graph Completion
Early explorations of LLM-driven KGC primarily focused on converting triples into natural language sequences to verify whether generative LLMs could address incompleteness effectively. FSKG [40] exemplifies this trend by introducing a generation–refinement pipeline that mitigates long-tail sparsity through staged generation. Building on this concept, MuKDC incorporated multi-stage knowledge distillation to generate coherent supplementary triples, further improving the coverage of long-tail relations. Extending beyond generation quality, KGs-LLM [130] demonstrated that zero-shot prompting, even without external knowledge, can iteratively extract graph components, thereby reducing annotation costs and scaling efficiently to domain-specific datasets. While these methods validated the feasibility of text-based KGC, KGLLM [131] further showed that such reformulation enables lightweight fine-tuning of smaller LLMs like LLaMA-7B, achieving competitive triple classification and relation prediction. Addressing the persistent challenge of long-tail entities, KICGPT [133] integrated a triple-based retriever with contextual prompting, directly encoding structural cues into LLM inputs. This approach avoids additional training while yielding strong few-shot performance, thus highlighting the efficiency of retrieval-augmented prompting in KGC.
To mitigate the structural information loss inherent in purely text-based approaches, researchers have developed prompt encoding strategies that embed richer graph context. RL-LLM [134] applied few-shot learning with multi-prompt optimization for relation prediction, showing that even minimal labeled data can yield competitive results in e-commerce KGs. Expanding the reasoning process, GoG [135] proposed a “think–search–generate” pipeline, enabling the synthesis of new triples without additional training, which strengthens reasoning over incomplete graphs. Complementing this, LLM-KGC [121] fused ontology and graph structure directly into prompts, aligning topological and semantic information to improve inductive reasoning. Further enhancing structure-aware inference, KoPA [136] introduced a knowledge prefix adapter to inject structural signals during LLM reasoning, thereby improving logical consistency in KGC predictions. These works collectively highlight a shift from simple text reformulation toward structurally enriched prompting, aiming to bridge the gap between LLM semantic priors and graph topology.
Building upon these prompting and structural integration advances, recent research has optimized LLM–graph fusion for few-shot KGC, entity alignment, and robustness in heterogeneous settings. MuKDC [141] extended its earlier generation framework by incorporating multimodal knowledge and consistency evaluation, boosting performance on long-tail few-shot tasks. In parallel, LLMKG [138] demonstrated that carefully engineered prompts allow both Mixtral-8x7b and GPT-4 to perform competitively in zero- and one-shot KGC across diverse metrics. Addressing entity alignment, DIFT [139] combined lightweight models with discriminative prompts to prevent alignment drift while improving overall KGC accuracy. In handling semantic ambiguity, CP-KGC [140] leveraged contextually adaptive prompts to disambiguate polysemous entities, particularly in quantized LLMs. Moving toward deeper structural integration, GLTW [128] fused Graph Transformers with LLMs to jointly capture local and global patterns, while GS-KGC [123] employed subgraph-based question answering with negative sampling to enhance missing triple identification. Together, these frameworks signal a transition from isolated prompt design to holistic, multimodal, and structure-aware LLM–graph systems.
@@
3.1.4 Evaluation for Handling Incompleteness in Graphs
We summarize the existing evaluation pipeline of LLM-based methods for incomplete graph learning, covering benchmark datasets, evaluation metrics, and downstream tasks (see Table 1).
Across the reviewed methods, commonly used benchmark datasets include Cora [81], Citeseer [85], PubMed [82], and ogbn-arxiv [83] for node classification tasks. These are citation networks where each node is a document with associated textual attributes. For large-scale knowledge graph tasks, datasets such as FB15k-237 [94], WN18RR [124], Wikidata5M [129], NELL [142], and UMLS [137] are widely adopted for link prediction and knowledge graph completion. Additionally, domain-specific or task-specific datasets are used to evaluate generalizability, including Amazon-Clothing and Amazon-Sports [100] for recommendation and link prediction, PeMS03/04/07 [88] for spatiotemporal forecasting, TQA/WTQ/TabFact [109, 112, 113] for table understanding, and GrailQA/WebQSP/KQA Pro [115, 116, 117] for knowledge graph question answering. Some works also utilize synthetic graphs [92, 135] to test reasoning capabilities in controlled environments.
A variety of evaluation metrics are used depending on the task. For classification, Accuracy, Macro-F1, and Micro-F1 are most common. Knowledge graph-related tasks report Mean Reciprocal Rank (MRR) and Hits@N as standard ranking metrics. Spatiotemporal forecasting relies on MAE, RMSE, and MAPE, while generative tasks such as question answering are evaluated using Exact Match, BLEU-4, and ROUGE-L.
The surveyed models address a wide range of downstream tasks, including node classification, link prediction, knowledge graph completion, anomaly detection, graph reasoning, question answering over knowledge graphs, and spatiotemporal forecasting. This diversity highlights the versatility of LLMs in handling incomplete graph data across domains and tasks.
@@ 3.2 Imbalance in Graphs
This section reviews the application of traditional methods and LLMs for learning from imbalanced graph data. The literature is categorized into two key research areas: (1) Class Imbalance and (2) Structure Imbalance. The relevant references and categorization are presented in Table 2.
Table 2: LLM-based methods for handling imbalance in graphs, grouped by domain and task, with representative methods, datasets, metrics, and downstream tasks.
Domain Tasks Method Typical Datasets Common Metrics Downstream Tasks Class Imbalance Graph Learning Node Classification LLM4NG [41] Cora [81], PubMed [82], ogbn-arxiv [83] Accuracy Few-shot Node Classification LLM-GNN [89] Cora [81], Citeseer [85], PubMed [82], ogbn-arxiv [83], ogbn-products [83], Wikics [90] Accuracy Label‐free Node Classification G2P2 [165] Cora [81], Art [166], Industrial [166] and Music Instruments [166] Accuracy, Macro-F1 Zero- and Few-shot Low-resource Text Classification LA-TAG [167] Cora [81], PubMed [82], Photo [168], Computer [168], and Children [168] Accuracy, Macro-F1 Zero- and Few-shot Low-resource Text Classification GSS-Net [169] Magazine Subscriptions [166], Appliances [166], Gift Cards [166] Accuracy, Precision, Recall, F1-score, MSE, RMSE, and MAE Sentiment Classification on Streaming E-commerce Reviews TAGrader [170] Cora [81], PubMed [82], ogbn-products [83], Arxiv-2023 [86] Accuracy Node Classification on Text-attributed Graphs SEGA [171] DAIC-WOZ [172], EATD [173] Macro-F1 Depression Detection SocioHyperNet [174] MBTI [174] Accuracy, AUC, Macro-F1, Micro-F1, IMP Examining Personality Traits Cella [175] Cora [81], Citeseer [85], PubMed [82], Wiki-cs [90] Accuracy, NMI, ARI, F1-score Label‐free Node Classification LLM-TIKG [176] threat-dataset [176] Precision, Recall, F1-Score Threat Intelligence Knowledge Graph Construction ANLM-assInNNER [177] NE dataset [177] Precision, Recall, F1-Score Robotic Fault Diagnosis Knowledge Graph Construction LLM-HetGDT [178] Twitter-HetDrug [178] Macro-F1, GMean Online Drug Trafficking Detection Prediction LLM-SBCL [179] biology [179], law [179], cardiff20102 [179], sydney19351 [179], and sydney23146 [179] Binary-F1, Micro-F1, Macro-F1, Accuracy Student Performance Prediction LKPNR [180] MIND [181] AUC, MRR, nDCG Personalized News Recommendation LLM-DDA [182] BCFR-dataset [182] AUC, AUPR, F1-score, Precision Computational Drug Repositioning Graph Completion KICGPT [133] FB15k-237 [94], WN18RR [124] MRR, Hits@N Link Completion KGCD [183] WN18RR [124], YAGO3-10 [94], WN18 [184] MRR, Hits@N Low-resource Knowledge Graph Completion Foundation Model GraphCLIP [185] ogbn-arXiv [83], Arxiv-2023 [86], PubMed [82], ogbn-products [83], Reddit [186], Cora [81], CiteSeer [85], Ele-Photo [168], Ele-Computers [168], Books-History [168], Wikics [90], Instagram [186] Accuracy Transfer Learning on Text-attributed Graphs Structure Imbalance Graph Learning Node Classification GraphEdit [187] Cora [81], Citeseer [85], PubMed [82] Accuracy Refining Graph Topologies Graph Completion SATKGC [188] WN18RR [124], FB15k-237 [94], Wikidata5M [189] MRR, Hits@N Knowledge Graph Completion MPIKGC [190] FB15k-237 [94], WN18RR [124], FB13 [132], WN11 [132] MR, MRR, Hits@N, Accuracy Knowledge Graph Completion LLM4RGNN [191] Cora [81], Citeseer [85], PubMed [82], ogbn-arxiv [83], ogbn-products [83] Accuracy Improving the Adversarial Robustness
@@ 3.2.1 Class-Imbalanced Graph Learning
Beyond data generation, another research direction emphasizes directly strengthening tail-class representations and rebalancing classifiers, often through retrieval or expert-based designs. For instance, RAHNet [198] combines a retrieval module with a second-stage classifier that applies max-norm and weight decay regularization. This joint design improves representation quality while reducing head-class dominance. Building on the idea of specialized modules, CoMe [199] introduces a collaborative-expert framework, where a general expert captures global shared patterns and a specific expert focuses on tail-class features; a collaboration regularizer further enforces complementarity, leading to more balanced performance. Extending this paradigm, KDEX [200] trains knowledge-diverse experts, each specializing in different structural or semantic patterns, and integrates their outputs via an expert allocation and fusion mechanism. Collectively, these approaches highlight how retrieval and expert-based architectures can complement generation strategies by directly enhancing the representation and robustness of tail classes.
@@
Another approach focuses on the contextual understanding and reasoning capabilities of LLMs, mining implicit relationships of minority classes through structure-aware reasoning. Both LLM-GNN [89] and Cella [175] proposed a zero-shot classification framework based on LLM semantic reasoning. By parsing node attributes and neighbor context, they directly generate category prediction confidence and get rid of the dependence on labeled data. G2P2 [165] introduces a graph-guided pretraining prompt framework, using LLMs to generate augmented text related to graph structure, enhancing the discriminability of low-resource classes in heterogeneous spaces. Meanwhile, LLM-TIKG [176] builds a threat intelligence knowledge graph, using LLMs to extract attack pattern association rules from unstructured text, and enhancing the representation of minority class threat entities through heterogeneous graph attention.
To improve generalization on sparse or long-tail categories in graphs, another line of work focuses on injecting external knowledge—either from open-domain corpora or domain-specific sources—into the learning process. KICGPT [133] enriches long-tail relational reasoning by dynamically constructing structure-aware prompts with contextual knowledge, significantly improving completion performance for sparse relations. In the recommendation domain, LKPNR [180] integrates user preference reasoning from LLMs and semantic path mining to address cold-start scenarios in long-tail news recommendation. KGCD [183] augments knowledge graph completion with pseudo-triplets generated from logical rules, guiding the model to focus on underrepresented relations. Other methods, such as [170], distill open-domain knowledge from LLMs into graph encoders through hierarchical knowledge transfer, while GraphCLIP [185] leverages cross-modal alignment to construct graph–text contrastive learning objectives for tail-category enhancement. In biomedical applications, LLM-DDA [182] infuses mechanistic knowledge into graph reconstruction, improving drug repurposing predictions for rare diseases. These works demonstrate that LLMs serve not just as predictors, but as external knowledge carriers that can guide graph learning systems beyond their native data distributions.